OpenStudy (anonymous):

The perimeter of a rectangle is 30 ft. The length is four times the width. find the length and width of the rectangle.

OpenStudy (anonymous):

30= 2(4x+x) 30= 8x+2x 30=10x x=3 width=3 lenght=12

Parth (parthkohli):

2(4x + x) = 30 2(5x) = 30 10x = 30 x = 3 So, LENGTH = 12 BREADTH = 3

Parth (parthkohli):

Perimeter of a rectangle = 2(a + b)

OpenStudy (saifoo.khan):

First step, suppose, width = x Length = 4x now you will understand. @yvone

OpenStudy (anonymous):

Thank you I needed a memory boost. exams blah!!!!

OpenStudy (saifoo.khan):

Please hit good answer. thanks.